OpenStudy (anonymous):

Find the area between the curves: y = x^3 - 9 x^2 + 14 x and y = -x^3 + 9 x^2 - 14 x

zepdrix (zepdrix):

Hi, \(\Large\bf \color{royalblue}{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (anonymous):

hi

zepdrix (zepdrix):

\[\Large\bf\sf y_1=x^3-9x^2+14x\]\[\Large\bf\sf y_2=-x^3+9x^2-14x\]Messing with y2 a little bit,\[\Large\bf\sf y_2=-(x^3-9x^2+14x) \qquad\implies\qquad y_2=-y_1\]

zepdrix (zepdrix):

So ummmm

zepdrix (zepdrix):

Do you understand how to find points of intersection? (This will tell us from where and to where we're integrating).

OpenStudy (anonymous):

yes make them equal to one another \

zepdrix (zepdrix):

Ok good. In doing so, you should get something like:\[\Large\bf\sf x^3-9x^2+14x\quad=\quad 0\]Which gives us 3 intersections: x=0, x=2, x=7

OpenStudy (anonymous):

right

zepdrix (zepdrix):

This one is a little tricky since we have 3 intersections. Here's the graph just to give you an idea of what we're dealing with: https://www.desmos.com/calculator/tzok9densw We'll have to setup 2 integrals. One from 0 to 2 and another from 2 to 7. And then add the integrals afterwards.

OpenStudy (anonymous):

i did!

OpenStudy (anonymous):

and i am still getting it wrong!

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits_0^2 y_1-y_2\;dx\quad+\quad\int\limits_2^7y_2-y_1\;dx\]

zepdrix (zepdrix):

So for our first integral:\[\Large\bf\sf \int\limits_0^2 x^3-9x^2+14x-(-x^3+9x^2-14x)\;dx\]

zepdrix (zepdrix):

\[\Large\bf\sf =2\int\limits_0^2 x^3-9x^2+14x\;dx\]

OpenStudy (anonymous):

shouldnt 14x be 28 x

OpenStudy (anonymous):

14x--14x=28 x right

OpenStudy (anonymous):

oh Nm..you divided by 2

zepdrix (zepdrix):

Yah I just noticed that we have 2 of everything :D factored a 2 out.

OpenStudy (anonymous):

right cool beans

OpenStudy (anonymous):

now we do the anti deriv. right

zepdrix (zepdrix):

Well that just takes care of the first region. We'll have to do the other integral as well. But ya let's go ahead and do the integration for this first one.

OpenStudy (anonymous):

which is \[((x^4)/4)-3x^3+7x^2\]

OpenStudy (anonymous):

did you get 16 as the answer for the first integral

zepdrix (zepdrix):

yes good.

zepdrix (zepdrix):

For the other integral, the order switches right?\[\Large\bf\sf \int\limits_2^7 (-x^3+9x^2-14x)-(x^3-9x^2+14x)\;dx\]

OpenStudy (anonymous):

right

zepdrix (zepdrix):

\[\Large\bf\sf =\quad -2\int\limits_2^7 x^3-9x^2+14x\;dx\]

OpenStudy (anonymous):

2588.25?

zepdrix (zepdrix):

bahh i dunno i didn't check it yet >.<

zepdrix (zepdrix):

Hmm it should work out to 187.5 I think.

OpenStudy (anonymous):

total?

zepdrix (zepdrix):

\[\Large\bf\sf -2\left[\frac{1}{4}x^4-3x^3+7x^2\right]_2^7\]

zepdrix (zepdrix):

Just for this piece, unless I punched something in wrong :o

OpenStudy (anonymous):

could be me ill do it again

OpenStudy (anonymous):

187.5?

OpenStudy (anonymous):

i got 203.5 total and i was right!

zepdrix (zepdrix):

Yayyy good job \c:/

zepdrix (zepdrix):

Tough problem cause of those two areas D:

OpenStudy (anonymous):

hey could i ask you antoher question or do i have to open up a new question thing again

OpenStudy (anonymous):

nm ill open another one and close this one thnx

zepdrix (zepdrix):

ya thats a good idea :D I'm kinda multitasking right now, so if i can't get to it it's good to have a fresh window so others can see it.